OpenStudy (anonymous):

which of the following values is NOT in the domain of the function h(x)=5-x/2x^2-8

OpenStudy (anonymous):

\[\frac{ 5 - x }{ 2x^2 - 8 } = \frac{ 1 }{ 2 }\times\frac{ 5-x }{ x^2 - 4 } = \frac{ 1 }{ 2 }\times\frac{ 5-x }{ (x+2)(x-2)} \] the denominator = 0 when x is -2 or 2. therefore they are not in the function's domain

OpenStudy (anonymous):

a.) 5 b.) 2.5 c.)-4 d.) -2

OpenStudy (anonymous):

thanks :)